### v1

@@ 4.2. Uniform estimates
$\mathopen{}\mathclose{{\left|\big\langle-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n}),\bm{\phi}\big\rangle}}\right|=\mathopen{}\mathclose{{\left|\int_{\Omega}(\bm{v}_{n}\cdot\nabla)\bm{v}_{n}\cdot\bm{\phi}\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}\|_{2^{\ast}}\,\|\nabla\bm{v}_{n}\|_{p}\,\|\bm{\phi}\|_{2^{\ast}_{s}},$
where $2^{\ast}=\frac{2d}{d-2}$ (for $d>2$, or arbitrarily large if $d=2$) and $2^{\ast}_{s}=\frac{2d}{d-2s}$ for $s<\frac{d}{2}$. Let us focus on the case $d>2$, once for $d=2$ the analysis is simpler. According to (4.2), $s>1+\frac{d}{2}$, and thus we can take $2^{\ast}_{s}=\infty$. Then, Hölder’s inequality requires
$\frac{d-2}{2d}+\frac{1}{p}<1\Leftrightarrow p>\frac{2d}{d+2}.$
The viscous term $\nu\,\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))$ is naturally controlled in $\mathrm{W}^{-1,p^{\prime}}(\Omega)^{d}$ and the same is true for the forcing term $\bm{f}$ in view of assumption (3.4).
Hence, a simultaneous control of all terms is achieved by taking $s=p$. As a result, for any dimension $d\geq 2$, it holds that
$\partial_{t}\bm{v}_{n}\in\mathrm{L}^{p^{\prime}}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d}).$ (4.17)
@@ 4.3. Passing to the limit as $n\to\infty$
Assuming $d>2$ (since the case $d=2$ is simpler), and combining (4.10) and (4.15) with the compact and continuous embeddings
$\mathrm{W}^{1,p}_{0}(\Omega)^{d}\hookrightarrow\hookrightarrow\mathrm{L}^{2}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r}(\Omega)^{d},\quad p>\frac{2d}{d+2},\quad r\leq\frac{2d}{d-2},\quad d\neq 2,$ (4.23)
we may apply Lemma 2.2-(2) with $s=0$, $p=2$, $p_{0}=2$, and $p_{1}=p$. This yields the existence of a function $\bm{v}$ such that, for some subsequence (again not relabeled),
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{in}\quad\mathrm{L}^{2}(Q_{T})^{d}.$ (4.24)
@@
Remark 4.4.
We can avoid the use of estimate (4.11), which requires the restriction $p\geq d$ when $d>3$ and is invoked in (4.14) for the application of the Aubin–Dubinskiĭ lemma (Lemma 2.2) to obtain the strong convergence (4.24). Instead, we can follow the argument outlined in Remark 4.3. Indeed, the uniform bounds provided by (4.10) and (4.17) are sufficient to apply the Aubin–Dubinskiĭ lemma directly, thereby guaranteeing the strong convergence of the Galerkin approximations.
@@ 4.5. Conclusion
We can see that (1)-(2) of Definition 3.1 are easily verified due to (4.18), (4.20) and (4.22). On the other hand, multiplying (4.4) by $\psi\in\mathrm{C}^{\infty}(0,T)$, with $\psi(T)=0$, integrating over $(0,T)$, using integration by parts in time, using (4.5), and the convergence results (4.7), (4.18), (4.21), (4.26) (observing here (3.8) for the choice of $\rho$ in (3.7)) and (4.36), we obtain (3.2) for any $\bm{\varphi}=\psi\bm{\psi}_{k}$, with $\bm{\psi}_{k}\in\mathbf{X}^{n}$, $\psi\in\mathrm{C}^{\infty}(0,T)$, and $\psi(T)=0$. Thus, by continuity and density, (3.2) holds true for any $\bm{\varphi}\in\mathbf{C}^{\infty}_{0,\operatorname{div}}(Q_{T})$. Finally, since $\partial_{t}\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$ and $\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$, we have $\bm{v}\in\mathrm{W}^{1,2}(0,T;\mathbf{V})\subset\mathrm{C}([0,T];\mathbf{V})$. Consequently, the initial condition $\bm{v}(0)=\bm{v}_{0}$, stated in (3) of Definition 3.1, is meaningful. This concludes the proof of Theorem 3.2 under assumption (4.1).
In the rest of this work, we focus on the most challenging case in the proof of Theorem 3.2, that is,
$\frac{2d}{d+2}<p<\frac{3d-4}{d}.$ (4.37)
This intermediate range of the power-law exponent $p$ is significantly more delicate because the embedding stated in (3.7) no longer holds. As a consequence, additional analytical tools are required to handle the nonlinear terms. In particular, it becomes necessary to perform a refined decomposition of the pressure into multiple components in order to compensate for the loss of compactness and to secure the weak convergence properties essential to the existence argument. The next section is devoted to developing the technical aspects of this pressure decomposition, which will play a crucial role in the construction of weak solutions for the case of $p$ satisfying (4.37).
@@ 5. Revisiting the pressure decomposition technique
6. Existence: small values of $p$
In this section, the proof of Theorem 3.2 when the power-law exponent satisfies (4.37) is carried out. We begin by regularizing problem (1.1) so that the results established in Section 4 remain applicable for the regularized problem (6.3) below. We choose a Lebesgue exponent $\beta$ such that
$\beta\geq\max\mathopen{}\mathclose{{\left\{\frac{3d-4}{d},p}}\right\}.$ (6.1)
@@ 6. Existence: small values of $p$
$\max\mathopen{}\mathclose{{\left\{\frac{3d-4}{d},p}}\right\}\leq\beta\leq d.$ (6.2)
@@ 6.1. Regularized problem
Since the exponent $\beta$ satisfies (6.1), it also meets condition (4.1). Therefore, the existence of a weak solution $\bm{v}_{n}$ to the problem (6.3), in the sense of Definition 6.1, can be directly inferred from the proof given in Section 4. In this case only an adaptation of assumption (3.3) is required,
@@ 6.2. A priori estimates
As we have seen in Section 4, and as $\beta$ satisfies (4.1), the space of test functions coincides with the space where the solutions to the problem (6.3) are constructed. In particular, the counterparts of (4.10) and (4.11) are also valid here in the same conditions of Lemma 4.1, using in addition the weakly lower semicontinuity of the norm,
@@ 6.3. Pressure recovery and decomposition
In the shear-thinning case ($1<p<2$), and in particular for $p$ satisfying (4.37), the reintroduction and decomposition of the pressure in the weak formulation constitute a structural necessity for establishing an existence result. This procedure enables the precise separation of the inertial, viscous, and convective contributions, as stated in Theorem 5.1. It further compensates for the lack of regularity of the velocity field, by providing the analytical framework required to control nonlinear terms under weaker integrability. In addition, the decomposition plays a key role in compactness and convergence arguments, allowing one to isolate suitable regular pressure components from more singular parts, thereby ensuring strong convergence of the nonlinear expressions in the weak formulation. Without this step, the interplay between the viscous nonlinearity and the limited compactness of the velocity field would obstruct the proof of existence. For that purpose, let us fix
@@
$\|\pi_{h,n}(t)\|_{2,r}\leq C_{1}\|\pi_{h,n}(t)\|_{2}\leq C_{2}\big(\|\bm{v}_{n}(t)-\bm{v}_{0,n}\|_{2}+\kappa\|\nabla\bm{v}_{n}(t)-\nabla\bm{v}_{0,n}\|_{2}\big),$ (6.28)
@@ 6.5. Passing to the limit $n\to\infty$
From above condition and (6.12), we deduce
@@ 7. Uniqueness
$\|\nabla\bm{w}(t)\|_{2}^{2}\leq\|\nabla\bm{w}(0)\|_{2}^{2}\exp(Ct),\ \text{ for all }\ \ t\in[0,T].$

### v2

@@ 4.2. Uniform estimates
$\begin{split}\mathopen{}\mathclose{{\left|\big\langle-\mathbf{div}\,(\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)),\bm{\phi}\big\rangle}}\right|=&\mathopen{}\mathclose{{\left|\int_{\Omega}\bm{v}_{n}(t)\otimes\bm{v}_{n}(t):\nabla\bm{\phi}\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}(t)\|_{2^{\ast}}^{2}\|\nabla\bm{\phi}\|_{\infty}\\ \leq&C\|\nabla\bm{v}_{n}(t)\|_{2}^{2}\|\bm{\phi}\|_{s,2},\end{split}$ (4.18)
for $s>1+\frac{d}{2}$. For the viscous term, we also have for any $\bm{\phi}\in\mathrm{W}_{0}^{s,2}(\Omega)^{d}$
$\begin{split}&\mathopen{}\mathclose{{\left|\big\langle\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n}(t))|^{p-2}\mathbf{D}(\bm{v}_{n}(t))),\bm{\phi}\big\rangle}}\right|\\ &=\mathopen{}\mathclose{{\left|\int_{\Omega}|\mathbf{D}(\bm{v}_{n}(t))|^{p-2}\mathbf{D}(\bm{v}_{n}(t))):\nabla\bm{\phi}\,\mathrm{d}x}}\right|\\ &\leq\|\mathbf{D}(\bm{v}_{n}(t))\|_{p-1}^{p-1}\|\nabla\bm{\phi}\|_{\infty}\leq C\|\mathbf{D}(\bm{v}_{n}(t))\|_{p}^{p-1}\|\bm{\phi}\|_{s,2},\end{split}$ (4.19)
for $s>1+\frac{d}{2}$ as well. As for the forcing term, we invoke assumption (3.4), in particular the fact that $\bm{f}\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-1,2}(\Omega)^{d})$. Moreover, the continuous embedding $\mathrm{W}^{-1,2}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-s,2}(\Omega)^{d}$, valid for every $s>1$, yields the same bound in the larger space. Consequently, we also have
$\bm{f}\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d}).$ (4.20)
Combining (4.17)–(4.20) with the estimates (4.12) and (4.13), we obtain (4.16).
We can now apply the Aubin-Dubinskiĭlemma with the compact embedding $\mathrm{W}^{1,2}_{0}(\Omega)^{d}\hookrightarrow\hookrightarrow\mathrm{L}^{2}(\Omega)^{d}$ and the continuous embedding $\mathrm{L}^{2}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-s,2}(\Omega)^{d}$, valid for every for $s>0$. Since, by (4.12), $\bm{v}_{n}$ is uniformly bounded in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})$ and, by (4.16), $\partial_{t}\bm{v}_{n}$ is uniformly bounded in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d})$, the Aubin-Dubinskiĭ lemma implies that, up to a subsequence,
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{ in }\ \mathrm{C}([0,T];\mathrm{L}^{2}(\Omega)^{d}).$ (4.21)
@@ 4.3. Passing to the limit as $n\to\infty$
On the other hand, (4.21) implies that, for some subsequence not relabeled,
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{in}\quad\mathrm{L}^{2}(0,T;\mathrm{L}^{2}(\Omega)^{d}).$ (4.27)
@@ 4.5. Conclusion
We can see that (1)-(2) of Definition 3.1 are easily verified due to (4.22), (4.24) and (4.26). On the other hand, multiplying (4.6) by $\psi\in\mathrm{C}^{\infty}(0,T)$, with $\psi(T)=0$, integrating over $(0,T)$, using integration by parts in time, using (4.7), and the convergence results (4.9), (4.22), (4.25), (4.29) (observing here (4.3) for the choice of $\rho$ in (4.1)) and (4.39), we obtain (3.2) for any $\bm{\varphi}=\psi\bm{\psi}_{k}$, with $\bm{\psi}_{k}\in\mathbf{X}^{n}$, $\psi\in\mathrm{C}^{\infty}(0,T)$, and $\psi(T)=0$. Thus, by continuity and density, (3.2) holds true for any $\bm{\varphi}\in\mathbf{C}^{\infty}_{0,\operatorname{div}}(Q_{T})$. Finally, since $\partial_{t}\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$ and $\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$, we have $\bm{v}\in\mathrm{W}^{1,2}(0,T;\mathbf{V})\subset\mathrm{C}([0,T];\mathbf{V})$. Consequently, the initial condition $\bm{v}(0)=\bm{v}_{0}$, stated in (3) of Definition 3.1, is meaningful. This concludes the proof of Theorem 3.2.
In the rest of this work, we focus on the most challenging case (see Theorem 3.3 and (3.8) therein):
$\frac{2d}{d+2}<p<\frac{d}{2}\qquad\mbox{and}\qquad d>3.$
This intermediate range of the power-law exponent $p$ is significantly more delicate in spatial dimensions $d>3$, because the embedding in (4.1) fails to hold in this setting. In lower dimensions, this embedding is valid, and the associated estimates can be carried out straightforwardly. However, for $d>3$, additional analytical tools are required to handle the nonlinear terms. In particular, it becomes necessary to perform a refined decomposition of the pressure into multiple components in order to compensate for the loss of compactness and to secure the weak convergence properties essential to the existence argument. The next section is devoted to developing the technical aspects of this pressure decomposition, which will play a crucial role in the construction of weak solutions for the case of $p$ satisfying (3.8).
@@ 5. Revisiting the pressure decomposition technique
6. Proof of Theorem 3.3
We begin by regularizing problem (1.1), a necessary step to ensure that Theorem 3.2 can be applied to the resulting problem presented in (6.3) below. We choose a Lebesgue exponent $\beta$ such that
$\beta\geq\frac{d}{2}\qquad\mbox{and}\qquad d>3.$ (6.1)
@@
$\frac{d}{2}\leq\beta\leq d\qquad\mbox{and}\qquad d>3.$ (6.2)
@@ 6.1. Regularized problem
Since the exponent $\beta$ satisfies (6.1), it also meets condition (3.6). Therefore, the existence of a weak solution $\bm{v}_{n}$ to the problem (6.3), in the sense of Definition 6.1, can be directly inferred from Theorem 3.2. In this case only an adaptation of assumption (3.3) is required,
@@ 6.2. A priori estimates
As we have seen in the proof Theorem 3.2 (see Section 4), and as $\beta$ satisfies (3.6), the space of test functions coincides with the space where the solutions to the problem (6.3) are constructed. In particular, the counterparts of (4.12) and (4.13) are also valid here in the same conditions of Lemma 4.1,
@@ 6.3. Pressure recovery and decomposition
In the particular case of $p$ satisfying (3.8), the reintroduction and decomposition of the pressure in the weak formulation constitute a structural necessity for establishing an existence result. This procedure enables the precise separation of the inertial, viscous, and convective contributions, as stated in Theorem 5.1. It further compensates for the lack of regularity of the velocity field, by providing the analytical framework required to control nonlinear terms under weaker integrability. In addition, the decomposition plays a key role in compactness and convergence arguments, allowing one to isolate suitable regular pressure components from more singular parts, thereby ensuring strong convergence of the nonlinear expressions in the weak formulation. Without this step, the interplay between the viscous nonlinearity and the limited compactness of the velocity field would obstruct the proof of existence. For that purpose, let us fix
@@
$\|\pi_{h,n}(t)\|_{2,r}\leq C_{1}\|\pi_{h,n}(t)\|_{2}\leq C_{2}\big(\|\bm{v}_{n}(t)-\bm{v}_{0,n}\|_{2}+\kappa\|\nabla\bm{v}_{n}(t)-\nabla\bm{v}_{0,n}\|_{2}\big),$
@@ 6.5. Passing to the limit $n\to\infty$
From the above condition and (6.12), we deduce
@@ 7. Uniqueness
$\|\nabla\bm{w}(t)\|_{2}^{2}\leq\|\nabla\bm{w}(0)\|_{2}^{2}e^{Ct},\ \text{ for all }\ \ t\in[0,T].$